# RIS-LAD: A Benchmark and Model for Referring Low-Altitude Drone Image Segmentation

### v1

@@ Abstract
Referring Image Segmentation (RIS), which aims to segment specific objects based on natural language descriptions, plays an essential role in vision-language understanding. Despite its progress in remote sensing applications, RIS in Low-Altitude Drone (LAD) scenarios remains underexplored. Existing datasets and methods are typically designed for high-altitude and static-view imagery. They struggle to handle the unique characteristics of LAD views, such as diverse viewpoints and high object density. To fill this gap, we present RIS-LAD, the first fine-grained RIS benchmark tailored for LAD scenarios. This dataset comprises 13,871 carefully annotated image-text-mask triplets collected from realistic drone footage, with a focus on small, cluttered, and multi-viewpoint scenes. It highlights new challenges absent in previous benchmarks, such as category drift caused by tiny objects and object drift under crowded same-class objects. To tackle these issues, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN). Rather than uniformly injecting all linguistic features, SAARN decomposes and routes semantic information to different stages of the network. Specifically, the Category-Dominated Linguistic Enhancement (CDLE) aligns visual features with object categories during early encoding, while the Adaptive Reasoning Fusion Module (ARFM) dynamically selects semantic cues across scales to improve reasoning in complex scenes. The experimental evaluation reveals that RIS-LAD presents substantial challenges to state-of-the-art RIS algorithms, and also demonstrates the effectiveness of our proposed model in addressing these challenges. The dataset and code will be publicly released soon at: https://github.com/AHideoKuzeA/RIS-LAD/.
Index Terms:
referring image segmentation, unmanned aerial vehicle, drone, aerial image analysis, benchmark dataset.
@@ I Introduction
Low-Altitude Drones (LAD), which typically operate below 200 meters, have become widely utilized in real-world applications due to their flexible deployment and high versatility [1, 2, 3]. This trend has sparked increasing research interest in vision tasks under LAD scenarios, leading to the development of benchmarks such as VisDrone [4] and UAV123 [5], which support object detection, tracking and other related tasks [6, 7, 8]. Recently, visual understanding in LAD scenarios [9, 10] has attracted increasing interest, especially in multi-modal tasks. Among them, Referring Image Segmentation (RIS) [11, 12, 13] is a fundamental perception task that aims to segment objects from images based on descriptions. Incorporating RIS into LAD systems enables LADs to better support a wider range of practical applications.
@@
However, existing RIS research primarily focuses on conventional scenes or high-altitude remote sensing imagery and lacks systematic investigations tailored to LAD imagery. Although previous Referring Remote Sensing Image Segmentation (RRSIS) studies [15, 14, 16] have provided valuable insights, they are mostly based on imagery taken from satellites or helicopters. As illustrated in Fig. 1(a), Referring Low-Altitude Drone Image Segmentation (RLADIS) presents several unique challenges compared to RRSIS: (1) diverse viewpoints resulting from lower and more flexible camera positions; (2) diverse illumination conditions; and (3) smaller and more densely distributed objects. These challenges introduce a substantial domain gap, which makes it difficult for existing RRSIS methods to generalize effectively to RLADIS tasks.
To bridge this gap, we propose RIS-LAD, the first dataset designed for fine-grained RLADIS. In addition, a semi-automatic annotation pipeline is also developed to ensure high-quality and efficient mask generation. The final dataset consists of 13871 image-text-mask triples and serves as a benchmark for evaluating RLADIS approaches. Furthermore, our analysis reveals two key challenges that hinder the direct adaptation of RRSIS methods to RLADIS: category drift and object drift. As illustrated in Fig. 1(b), category drift arises when the object occupies only a small region of the image, leading the model to mistakenly focus on larger, semantically similar objects. Object drift is caused by a high density of instances in the same category, which makes it difficult for the model to correctly identify the object described by the expression.
Motivated by the issues of category and object drift in RLADIS, we introduce the Semantic-Aware Adaptive Reasoning Network (SAARN), which decouples linguistic features and injects them at appropriate stages of the network to ensure semantic alignment. In contrast to existing decoupling-based approaches that inject all linguistic components uniformly across modules [17, 18], SAARN follows a more targeted strategy. Specifically, the Category-Dominated Linguistic Enhancement (CDLE) module injects only class-level linguistic features in the encoder stage. This enables early visual representations to be aligned with accurate category semantics. Meanwhile, the global linguistic features are selectively integrated via class-guided gating, which further reinforces semantic consistency. In the multi-scale fusion stage, the Adaptive Reasoning Fusion Module (ARFM) performs scale-aware enhancement by dynamically weighting semantic cues across multi-scale features. This mechanism simulates a reasoning process that guides the model to infer the most semantically consistent instance among densely packed objects of the same category.
@@
  • We introduce RIS-LAD, the first Referring Image Segmentation (RIS) dataset tailored for Low-Altitude Drone (LAD) scenarios, addressing the lack of fine-grained annotated data in this domain. It features diverse LAD scenes with rich textual descriptions and high-quality segmentation masks.
  • To address category drift and object drift in RLADIS, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN), which emphasizes semantic category alignment and selectively integrates the most relevant linguistic features at appropriate feature scales.
  • Through comprehensive experiments, we investigate the challenges in RLADIS and construct a benchmark tailored for this task. Meanwhile, the proposed SAARN achieves state-of-the-art performance on the RIS-LAD benchmark.
@@ II-1 Referring Image Segmentation (RIS)
Referring Image Segmentation (RIS) aims to achieve pixel-level localization of objects in images based on natural language expressions [19]. To address key limitations in RIS, recent research has explored advanced strategies. IteRPrimE [12] refines visual activation maps through an iterative Grad-CAM refinement strategy and adopts a primary-word emphasis module to improve the identification of key semantic components in language expressions. VATEX [20] leverages a CLIP-based prior module to generate heatmaps, which enforces semantic consistency and improve contextual alignment between language and vision. ASDA [21] utilizes adaptive selection and dual-alignment methods to enable adaptive alignment of visual and linguistic features. DeRIS [22] decomposes the RIS task into perception and cognition components, and employs a loop-back synergy mechanism to achieve bidirectional enhancement between the two modules. However, while these approaches perform well on conventional images, their effectiveness significantly declines when applied to the LAD imagery, due to substantial domain gaps between the two types of imagery.
@@ II-2 Referring Remote Sensing Image Segmentation (RRSIS)
RRSIS is a domain-specific extension of the RIS task for remote sensing images, first introduced by [23]. Existing RRSIS benchmarks [24, 25, 26] are mainly constructed from high-altitude sources such as Google Earth and GF-2. Recent methods [27, 28] have advanced the performance of RRSIS through different strategies. FIANet [17] disentangles various components of the referring expression and incorporates a text-aware multi-scale enhancement module to improve multi-modal discrimination. RMSIN [25] designs rotation-aware modules to precisely segment objects based on the unique properties of remote sensing imagery. RSRefSeg integrates a CLIP-based [29] module to capture implicit visual activations and uses them as guidance prompts for SAM [30]. DiffRIS [31] leverages the semantic understanding capabilities of pretrained text-to-image diffusion models to improve cross-modal alignment. Although some benchmarks have considered low-altitude scenarios (Li and Zhao 2025), drone imagery still presents unique challenges. These include tiny object sizes and high object density, which result from low flying altitudes and diverse viewpoints, significantly complicating both semantic grounding and visual recognition.
@@ III RIS-LAD Dataset
This section presents the construction pipeline of the proposed RIS-LAD dataset in detail.
III-1 Image Selection
@@
RIS-suitable images are selected from the publicly available CODrone dataset [32], which is designed for LAD-oriented object detection. The selected images cover diverse real-world urban scenes under different lighting conditions and from various viewpoints. Following common procedures in constructing RRSIS datasets [26, 24, 33], the selected high-resolution images are divided into uniform $\times 1080$ 1080 tiles, thereby preserving fine-grained details. A manual screening step is conducted to ensure that the retained tiles exhibit representative characteristics of LAD imagery.
III-2 Object Selection
[…]
III-3 Mask and Expression Generation Pipeline
The Segment Anything Model 2 (SAM-2) [34] is utilized to generate high-quality instance masks in an automated manner. Specifically, oriented bounding boxes provided by the CODrone serve as precise prompts to guide the segmentation process.
As illustrated in Fig. 2, referring expressions are generated using the multi-modal large language model (MM-LLM) Qwen2.5-VL [35]. In order to capture both instance-level appearance and spatial attributes, each object is cropped and upsampled to improve visual clarity. The cropped instance and the corresponding full image, with a red bounding box serving as the location marker, are then fed into Qwen2.5. This dual-input strategy helps mitigate the challenges posed by small object sizes. These objects often hinder the MM-LLM’s ability to capture fine-grained visual features and generate accurate spatial descriptions when relying solely on raw coordinates. As shown in Fig. 3, we further visualize linguistic diversity through word clouds generated from the refined referring expressions. The complete prompting strategy is provided in the supplementary materials.
Figure 3: Word cloud for referring expressions in the RIS-LAD dataset
III-4 Manual Refinement and Dataset Splitting
All automatically generated annotations are carefully verified and refined to ensure strong semantic alignment among images, referring expressions, and masks. This process yields a total of 13,871 high-quality image-text-mask triplets. These samples are then split into training, validation, and test sets with a 7:1:2 ratio.
Dataset Image Source Shooting Angle Nighttime Scene RefDIOR [24] Google Earth fixed NWPU-Refer [33] Google Earth fixed RISBench [24] Google Earth, GF-2, JL-1 fixed RefSegRS [26] Helicopter (above 1000m) fixed RRSIS-D [25] Google Earth fixed RIS-LAD Low-Altitude Drone $30^{\circ}$$60^{\circ}$
@@ III-4 Manual Refinement and Dataset Splitting
TABLE I: Comparison of RIS-LAD and existing referring remote sensing image segmentation datasets.
@@
Figure 4: Illustration of the proposed Semantic-Aware Adaptive Reasoning Network (SAARN). The framework comprises two core components: Category-Dominated Linguistic Enhancement (CDLE, top) and Adaptive Reasoning Fusion Module (ARFM, bottom). CDLE injects class-level linguistic features at the encoder stage to mitigate early-stage semantic misalignments. ARFM adaptively integrates multi-scale and linguistic features to reason about the correct object in scenes with dense object distributions. $\omega$ and $\varphi$ denote linear and nonlinear mappings, respectively.
III-5 Comparison of RIS-LAD and RRSIS datasets.
@@
As summarized in Table I, RIS-LAD exhibits distinct characteristics that distinguish it from existing RRSIS datasets. First, in contrast to RRSIS datasets that rely on fixed top-down viewpoints, RIS-LAD contains imagery captured from low-altitude drones at oblique angles between $30^{\circ}$ and $60^{\circ}$, thus offering greater variation in perspective and object appearance. Secondly, RIS-LAD uniquely includes nighttime scenes, enabling evaluation under challenging illumination conditions. Furthermore, it employs a semi-automatic annotation pipeline that integrates MM-LLMs with manual verification, thereby ensuring both annotation quality and efficiency.
@@ IV-A Overview
As illustrated in Fig. 4, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN), a framework designed to address category drift and object drift in Referring Low-Altitude Drone Image Segmentation (RLADIS). Built on RMSIN, SAARN incorporates two key components: the Category-Dominated Linguistic Enhancement (CDLE) and the Adaptive Reasoning Fusion Module (ARFM).
Given an input image, a Swin Transformer [36] encoder extracts visual features, while the referring expression is decomposed into global, class-level, and descriptive components, yielding three types of linguistic features: $l$, $c$, and $d$. For instance, in the expression ”The red car in the middle of the image parked in front of the white car,” the entire sentence corresponds to the global linguistic component; ”car” represents the class-level component; and the phrase ”red [masked] in the middle of the image parked in front of the white car,” excluding the category word, serves as the descriptive component.
At the encoder stage, CDLE injects class-level guidance through $c$ to align early visual representations with the correct category. Subsequently, it integrates $l$ to construct category-aware visual semantics. In the multi-scale fusion stage, ARFM employs an adaptive integration of pyramid fooled features with $l$, $c$, and $d$ across different scales. It assigns adaptive weights based on semantic alignment and spatial granularity, allowing each scale to emphasize the most relevant linguistic cues. The final fused representation is then passed to a mask decoder to generate the segmentation output.
@@ IV-B Category-Dominated Linguistic Enhancement
In order to address category drift, we propose the Category-Dominated Linguistic Enhancement (CDLE) module, which introduces only class-level linguistic features to align early visual representations with accurate semantic categories. Descriptive linguistic components are deliberately excluded, as they may align with visually similar but incorrect objects, thereby worsening category drift. This module is integrated into each stage of the visual encoder to enable fine-grained cross-modal interaction.
@@
where each $\omega$ denotes a convolutional layer applied along the channel dimension, and $d$ is the scaling factor.
Although the category-guided attention $\alpha^{c}$ effectively captures semantic alignment between the input and class-level linguistic cues, it tends to over-emphasize category semantics while suppressing other informative visual contexts. To offset this, a residual gate mechanism is employed, which facilitates adaptive weighting between $\alpha^{c}$ and the original feature $x$:
@@
This design ensures that $l$ is selectively activated in regions corresponding to $c$, thus providing a safeguard mechanism for the fusion between $l$ and $f^{c}$. The output $f^{l}$ retains the discriminative semantics of the referring expression while maintaining strong alignment with the correct category. This effectively mitigates attention drift toward visually or semantically similar but incorrect instances.
@@
V Adaptive Reasoning Fusion Module
The Adaptive Reasoning Fusion Module (ARFM) has been proposed to address the problem of object drift induced by densely packed same-category instances in LAD scenarios. Inspired by the human perception process, in which reasoning typically begins with coarse spatial localization and gradually refines toward specific object attributes, ARFM adopts a progressive fusion strategy. It establishes a dynamic weighting scheme between the fused multi-scale features $X^{f}=\{x^{f}_{1},x^{f}_{2},x^{f}_{3},x^{f}_{4}\}$ and the linguistic features $l$, $c$, and $d$. Thus, the model can capture scale-specific semantic cues more effectively, enabling semantic-aware attention across spatial resolutions. For instance, fine-grained details, such as color or shape, can be better captured at higher spatial resolutions, while coarse spatial descriptors such as “top-left corner,” are more salient at lower resolutions.
$X^{f}$ is first aligned using pyramid pooling, which is then downsampled to a shared resolution and channel dimension, resulting in $X^{fp}\in\mathbb{R}^{B\times C^{\prime}\times H^{\prime}\times W^{\prime}}$. These features are then fed into the Adaptive Reasoning Fusion Block (ARFB), which performs multi-branch, cross-modal reasoning guided by linguistic features. As illustrated in Fig. 5, ARFB contains three parallel linguistic branches that attend to $c$, $d$ and $l$, respectively. $X^{fp}$ is projected into a shared embedding space and interacts with each linguistic feature. Formally, for each linguistic feature $s\in\{l,d,c\}$, we compute:

### v2

@@ Abstract
Referring Image Segmentation (RIS), which aims to segment specific objects based on natural language descriptions, plays an essential role in vision-language understanding. Despite its progress in remote sensing applications, RIS under Low-Altitude Drone (LAD) scenarios remains underexplored, as existing datasets and methods are typically designed for high-altitude and static-view imagery. They struggled to handle the unique characteristics of LAD views, such as diverse viewpoints and high object density. In this paper, we propose RIS-LAD, the first fine-grained RIS benchmark tailored for LAD scenarios, featuring 13,871 meticulously annotated image-text-mask triplets collected from real-world drone footage with emphasis on small, densely cluttered objects and multi-view perspectives. It highlights new challenges not present in previous benchmarks, such as category drift caused by tiny objects and object drift under crowded objects of the same class. Additionally, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN), which decomposes and adaptively routes semantic information to different network stages rather than uniformly injecting all linguistic features. Specifically, the Category-Dominated Linguistic Enhancement (CDLE) aligns visual features with object categories during early encoding, while the Adaptive Reasoning Fusion Module (ARFM) dynamically selects semantic cues across scales to enhance reasoning in complex scenes. Extensive experiments reveal that RIS-LAD presents substantial challenges to state-of-the-art RIS algorithms, and also demonstrate the effectiveness of our proposed model in addressing these challenges. The dataset and code will be publicly released soon at: https://github.com/AHideoKuzeA/RIS-LAD-A-Benchmark-and-Model-for-Referring-Low-Altitude-Drone-Image-Segmentation.
@@ I Introduction
L, which typically operate below 200 meters, have become widely utilized in real-world applications due to their flexible deployment and high versatility [1, 2, 3]. This trend has sparked increasing research interest in vision tasks under LAD scenarios, leading to the development of benchmarks such as VisDrone [4] and UAV123 [5], which support object detection, tracking and other related tasks [6, 7, 8]. Recently, visual understanding in LAD scenarios [9, 10] has attracted growing interest, especially in multi-modal tasks. Among them, Referring Image Segmentation (RIS) [11, 12, 13] is a fundamental perception task that aims to segment objects from images based on descriptions. Incorporating RIS into LAD systems enables LADs to better accommodate a wider range of practical applications.
@@
However, existing RIS research mainly focuses on conventional scenes or high-altitude remote sensing imagery, leaving the distinctive properties of LAD imagery largely unexplored in the current literature. Although previous Referring Remote Sensing Image Segmentation (RRSIS) studies [15, 14, 16] have provided valuable insights, they are mostly based on imagery taken from satellites or helicopters. As shown in Fig. 1(a), Referring Low-Altitude Drone Image Segmentation (RLADIS) presents several unique challenges compared to RRSIS: (1) diverse viewpoints due to lower and more flexible camera positions; (2) diverse illumination conditions; and (3) smaller and more densely distributed objects. These challenges introduce a substantial domain gap that hinders the effective generalization of current RRSIS methods to RLADIS tasks.
To bridge this gap, we propose RIS-LAD, the first dataset designed for fine-grained RLADIS, accompanied by a semi-automatic annotation pipeline that ensures high-quality referring expression generation and annotation efficiency. The final dataset contains 13,871 image-text-mask triples, providing a benchmark for evaluating RLADIS methods. Furthermore, our analysis reveals two key challenges that hinder the direct adaptation of RRSIS methods to RLADIS: category drift and object drift. As shown in Fig. 1(b), category drift arises when the object occupies only a small region of the image, misleading the model to focus on larger, semantically similar objects. Object drift, on the other hand, results from a high density of instances belonging to the same category, making it difficult for the model to correctly identify the specific object referenced in the expression.
Motivated by these drift issues, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN), which decouples linguistic features and injects them at appropriate stages of the network to ensure semantic alignment. Unlike existing decoupling-based approaches that uniformly inject all linguistic components across modules [17, 18], SAARN employs a more targeted strategy. Specifically, the Category-Dominated Linguistic Enhancement (CDLE) module injects only class-level linguistic features into the encoder. This injection aligns early visual representations with accurate category-level semantics. Meanwhile, global linguistic features are selectively integrated via class-guided gating, reinforcing semantic consistency. In the multi-scale fusion stage, the Adaptive Reasoning Fusion Module (ARFM) performs scale-aware enhancement by dynamically weighting semantic cues across multi-scale features. This mechanism simulates a reasoning process, guiding the model to infer the most semantically consistent instance among densely packed objects of the same category.
@@
  • We propose RIS-LAD, the first fine-grained referring image segmentation dataset tailored for low-altitude drone scenarios. It offers diverse LAD scenes with rich textual descriptions and high-quality segmentation masks.
  • To address category drift and object drift in RLADIS, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN). This model focuses on aligning semantic categories and selectively integrates the most relevant linguistic features at optimal feature scales.
  • We conducted extensive experiments to investigate the challenges inherent in RLADIS and construct a dedicated benchmark for this task. Our proposed SAARN achieves state-of-the-art performance on the RIS-LAD benchmark.
@@ II-1 Referring Image Segmentation (RIS)
aims to localize target objects at the pixel level in an image based on natural language expressions [19]. A representative line of methods focuses primarily on enhancing vision-language alignment to improve segmentation performance [20, 21, 22]. VATEX [23] leverages a CLIP-based prior module to generate heatmaps, which enforces semantic consistency and improves contextual alignment between language and vision. ASDA [24] employs a dual-alignment mechanism with dynamic feature selection to better align visual and linguistic modalities. Other approaches explore novel architectural designs for RIS [25, 26, 27]. IteRPrimE [12] refines visual activation maps through an iterative Grad-CAM refinement strategy and adopts a primary-word emphasis module to improve the identification of key semantic components in language expressions. Although these approaches demonstrate promising performance on conventional imagery, their effectiveness deteriorates notably when applied to LAD imagery.
@@ II-2 Referring Remote Sensing Image Segmentation (RRSIS)
is a domain-specific extension of the RIS task for remote sensing images, first introduced by [28]. Given the high resolution and small object sizes in remote sensing images, many recent approaches have adopted multi-scale feature fusion [29, 30]. FIANet [17] disentangles various components of the referring expression and incorporates a text-aware multi-scale enhancement module to improve multi-modal discrimination. RMSIN [31] designs rotation-aware modules to precisely segment objects based on the unique properties of remote sensing imagery. Meanwhile, pre-trained large vision models have drawn growing interest in this field [32]. RSRefSeg integrates a CLIP-based [33] module to capture implicit visual activations and uses them as guidance prompts for SAM [34]. For benchmarks, existing RRSIS datasets [35, 31, 36] are mainly constructed from high-altitude sources such as Google Earth and GF-2, which feature scenes with larger objects and fixed viewpoints. Although some studies have explored low-altitude scenarios [9], challenges such as tiny objects and high instance density persist due to diverse viewpoints. Existing LAD-based datasets like UAVid-RIS and VDD-RIS [10] remain limited in object categories and expression granularity, restricting their ability to support fine-grained RIS tasks.
@@ III RIS-LAD Dataset
This section introduces the construction process and key characteristics of the RIS-LAD. Full dataset details and ethical statements are provided in the appendix.
III-1 Image and Category Selection
@@
The suitable images for the RIS task are selected from the publicly available CODrone dataset [37], which is designed for LAD-oriented object detection. The selected images encompass diverse real-world urban scenes under various lighting conditions and from different viewpoints. Following common procedures for constructing RRSIS datasets [36, 35, 38], the images are divided into uniform $1080\times 1080$ patches, thus preserving fine-grained details. A manual screening step is performed to ensure that the retained tiles exhibit representative characteristics of the LAD imagery.
Following mainstream LAD object detection benchmarks [4, 8, 5], we select eight representative object categories: people, car, motor, bicycle, tricycle, truck, bus and boat. These categories are commonly encountered and crucial for various drone-based applications. To better reflect real-world challenges, we apply a size constraint to ensure that more than $90\%$ of the annotated instances possess a mask coverage ratio below 0.1. This constraint highlights the prevalence of small and visually inconspicuous objects in LAD scenarios.
Figure 3: Word cloud for referring expressions in the RIS-LAD dataset.
Dataset Image Source Shooting Angle Nighttime Scene
RefDIOR Google Earth fixed
NWPU-Refer Google Earth fixed
RISBench Google Earth, GF-2, JL-1 fixed
RefSegRS Helicopter (above 1000m) fixed
RRSIS-D Google Earth fixed
RIS-LAD Drone (30m-100m) $30^{\circ}$$60^{\circ}$
@@
TABLE I: Comparison of RIS-LAD and publicly available referring remote sensing image segmentation datasets.
III-2 Mask and Expression Generation Pipeline
Segment Anything Model 2 (SAM-2) [39] is employed to automatically generate high-quality instance masks, where the oriented bounding boxes from CODrone provide precise prompts to guide the segmentation process.
As shown in Fig. 2, referring expressions are generated using the multi-modal large language model (MM-LLM) Qwen2.5-VL [40]. To capture both instance-level appearance and spatial attributes, each object is cropped and upsampled to improve visual clarity. The cropped instance and the corresponding full image, marked with a red bounding box that indicates the location, are then fed into Qwen2.5-VL. This dual-input strategy addresses challenges associated with small object sizes by the MM-LLM’s ability to capture fine-grained visual features and generate accurate spatial descriptions when relying solely on raw coordinates. As shown in Fig. 3, we further visualize the linguistic diversity through word clouds generated from the refined referring expressions.
III-3 Manual Refinement and Dataset Splitting
All automatically generated annotations are carefully verified and refined to ensure strong semantic alignment among images, referring expressions, and masks. This process yields a total of 13,871 high-quality image-text-mask triplets, randomly divided into training, validation, and test sets with a 7: 1: 2 ratio.
@@ III-3 Manual Refinement and Dataset Splitting
Figure 4: Illustration of the proposed Semantic-Aware Adaptive Reasoning Network (SAARN). The framework comprises two core components: Category-Dominated Linguistic Enhancement (CDLE, top) and Adaptive Reasoning Fusion Module (ARFM, bottom). To mitigate early-stage semantic misalignments, CDLE injects class-level linguistic features at the encoder stage, while ARFM adaptively integrates multi-scale and linguistic features to reason about the correct object in scenes with dense object distributions. $\omega$ and $\varphi$ denote linear and nonlinear mappings, respectively.
III-4 Comparison of RIS-LAD and RRSIS datasets.
@@
As summarized in Table I, RIS-LAD exhibits distinct characteristics that distinguish it from existing RRSIS datasets [35, 38, 36, 31, 30]. First, in contrast to RRSIS datasets that rely on fixed top-down viewpoints, RIS-LAD contains imagery captured from low-altitude drones at oblique angles between $30^{\circ}$ and $60^{\circ}$, thus offering greater variation in perspective and object appearance. Secondly, RIS-LAD uniquely includes nighttime scenes, enabling evaluation under challenging illumination conditions. Furthermore, it employs a semi-automatic annotation pipeline that integrates MM-LLMs with manual verification, ensuring both annotation quality and efficiency.
@@ IV-A Overview
As shown in Fig. 4, we propose the Semantic-Aware Adaptive Reasoning Network (SAARN), a framework designed to address category drift and object drift in RLADIS. SAARN incorporates two key components: the Category-Dominated Linguistic Enhancement (CDLE) and the Adaptive Reasoning Fusion Module (ARFM).
Given an input image and its corresponding referring expression, a Swin Transformer [41] encoder extracts visual features, and the expression is decomposed into global, class-level, and descriptive components, yielding three types of linguistic features: $l$, $c$, and $d$. As shown in the bottom-left of Fig. 4, these linguistic features correspond to the complete expression, the category name, and the descriptive content excluding the category term, respectively. Such disentanglement facilitates targeted modeling of semantic intent, object class information, and detailed spatial or attribute cues, which aligns with the intuitive human reasoning process of first locating the described region or identifying all category-matching objects, and then selecting the target based on specific attributes.
At the encoder stage, CDLE injects class-level guidance through $c$ to align early visual representations with the correct category. Subsequently, it integrates $l$ to construct category-dominated fused representations, which are then combined via residual fusion with output from the previous layer and the Visual Enhancement module [31] for multi-scale fusion. In the multi-scale fusion stage, ARFM employs an adaptive integration of pyramid pooled features with $l$, $c$, and $d$ across different scales. It assigns adaptive weights based on semantic alignment and spatial resolutions, allowing features at each scale to emphasize the most relevant linguistic cues. The final fused representation is then passed to a mask decoder to generate the segmentation output.
@@ IV-B Category-Dominated Linguistic Enhancement
In order to address category drift, we propose the Category-Dominated Linguistic Enhancement (CDLE) module, which selectively injects class-level linguistic features to precisely align early visual representations with the correct categories. Descriptive linguistic components are deliberately excluded to prevent alignment with visually similar but incorrect objects, thus mitigating category drift. This module is integrated into each stage of the visual encoder to enable fine-grained cross-modal interaction.
@@
where each $\omega$ denotes a $1\times 1$ convolutional layer applied along the channel dimension, and $d$ is the scaling factor.
Although category-guided attention $\alpha^{c}$ effectively captures semantic alignment between input and class-level linguistic cues, it tends to over-emphasize category semantics while suppressing other informative visual contexts. To offset this, a residual gate mechanism is employed, which facilitates adaptive weighting between $\alpha^{c}$ and the original feature $x$:
@@
This design ensures that $l$ is selectively activated in regions corresponding to $c$, thus providing a safeguard mechanism for the fusion between $l$ and $f^{c}$. The output $f^{l}$ retains the discriminative semantics of the referring expression while maintaining a strong alignment with the correct category. This effectively mitigates attention drift toward visually or semantically similar but incorrect instances.
@@
IV-C Adaptive Reasoning Fusion Module
The Adaptive Reasoning Fusion Module (ARFM) has been proposed to address the problem of object drift induced by densely packed same-category instances in LAD scenarios. Inspired by the human perception process, in which reasoning typically begins with coarse spatial localization and gradually refines toward specific object attributes, ARFM adopts a progressive fusion strategy. It establishes a dynamic weighting scheme between the fused multi-scale features $X^{f}=\{x^{f}_{1},x^{f}_{2},x^{f}_{3},x^{f}_{4}\}$ and the linguistic features $l$, $c$, and $d$. Thus, the model can capture scale-specific semantic cues more effectively, enabling semantic-aware attention across spatial resolutions. Fine-grained attributes like color and shape are better preserved at higher resolutions, whereas coarse spatial cues such as “top-left corner” become more distinguishable at lower resolutions.
$X^{f}$ is first aligned using pyramid pooling, which is then downsampled to a shared resolution and channel dimension, resulting in $X^{fp}\in\mathbb{R}^{B\times C^{\prime}\times H^{\prime}\times W^{\prime}}$. These features are then fed into the Adaptive Reasoning Fusion Block (ARFB), which performs cross-modal reasoning across multiple branches, guided by linguistic features. As shown in Fig. 5, ARFB contains three parallel linguistic branches that attend to $c$, $d$ and $l$, respectively. $X^{fp}$ is projected into a shared embedding space and interacts with each linguistic feature. Formally, for each linguistic feature $s\in\{l,d,c\}$, we compute: